# Quantum router of silicon-vacancy centers via a diamond waveguide

### v1

@@ IV Discussion and conclusion
Our solid-state quantum router is realizable under state-of-the-art nanofabrication techniques, including advanced strain control and precision ion implantation. It maintains complete compatibility with current experimental platforms [80, 81]. As demonstrated in Ref. [54], three arrays of SiV centers are successfully integrated into a diamond cantilever at precisely designed positions. In general, the proposed setup can be experimentally realized by embedding an array of SiV centers into a one-dimensional diamond waveguide. The spacing between SiV centers embedded in diamond waveguides in this paper can reach 20 $\mathrm{nm}$, which is significantly smaller than the phonon wavelength. Nevertheless, recent studies have demonstrated that strong phonon correlations can indeed be established and harnessed for efficient heat transport even in nanostructures with periodic features smaller than the phonon wavelength [82, 74]. Moreover, the spatial precision in fabricating SiV centers has been reported at the sub-30 $\mathrm{nm}$ level [83], further supporting the feasibility of achieving such small inter-center distances in diamond waveguides. Although achieving precise positional control of SiV centers and fabricating waveguides with cross-sectional areas as small as 100 $\mathrm{nm}^{2}$ remains challenging with current technologies, continued progress in diamond nanostructure fabrication is expected to address these limitations. The diamond waveguide, with material parameters $\varrho=3500\,\text{kg/m}^{3}$, $E=1050\,\text{GPa}$, and $\nu=0.2$, supports phonon propagation at a velocity $v\approx 1\times 10^{4}\,\text{m/s}$. These satisfy the requirement of our scheme.
In summary, we have proposed a quantum router based on SiV centers coupled to phonon modes in a one-dimensional diamond waveguide. We have discovered that the emergence of bound states in the energy spectrum of the total system enables persistent quantum entanglement and facilitates noise-resilient state transfer among distant SiV centers. This mechanism allows the realization of a multi-node quantum router without the need for complex external controls. By harnessing the bound-state mechanism, we have designed a scalable non-Markovian quantum router that simultaneously solves the challenges of distance limitations, fidelity degradation, and parallel interconnections faced by conventional quantum routers. Our finding establishes an engineered solid-state system as a viable platform for quantum routers and opens an avenue for developing quantum networks.
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (Grants No. 12275109 and No. 12247101), the Innovation Program for Quantum Science and Technology of China (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), and the Natural Science Foundation of Gansu Province (No. 22JR5RA389).

### v2

@@ IV Discussion and conclusion
Our solid-state quantum router is realizable under state-of-the-art nanofabrication techniques, including advanced strain control and precision ion implantation. It maintains complete compatibility with current experimental platforms [72, 73]. As demonstrated in Ref. [43], three arrays of SiV centers are successfully integrated into a diamond cantilever at precisely designed positions. In general, the proposed setup can be experimentally realized by embedding an array of SiV centers into a one-dimensional diamond waveguide. The spacing between SiV centers embedded in diamond waveguides in this paper can reach 20 $\mathrm{nm}$, which is significantly smaller than the phonon wavelength. Nevertheless, recent studies have demonstrated that strong phonon correlations can indeed be established and harnessed for efficient heat transport even in nanostructures with periodic features smaller than the phonon wavelength [74, 66]. Moreover, the spatial precision in fabricating SiV centers has been reported at the sub-30-$\mathrm{nm}$ level [75], further supporting the feasibility of achieving such small intercenter distances in diamond waveguides. Although achieving precise positional control of SiV centers and fabricating waveguides with cross-sectional areas as small as 100 $\mathrm{nm}^{2}$ remains challenging with current technologies, continued progress in diamond nanostructure fabrication is expected to address these limitations. The diamond waveguide, with material parameters $\varrho=3500\,\text{kg/m}^{3}$, $E=1050\,\text{GPa}$, and $\nu=0.2$, supports phonon propagation at a velocity $v\approx 1\times 10^{4}\,\text{m/s}$. These satisfy the requirement of our scheme. As a final remark, our scheme also works if the initial quantum information is encoded in another SiV center instead of the first one. This inherent flexibility is another advantage of our router architecture: any node in the network can act as the input port, enabling dynamic and reconfigurable quantum information routing based on need, rather than being limited to a fixed and pre-determined path.
In summary, we have proposed a quantum router based on SiV centers coupled to phonon modes in a one-dimensional diamond waveguide. We have discovered that the emergence of bound states in the energy spectrum of the total system enables persistent quantum entanglement and facilitates noise-resilient state transfer among distant SiV centers. This mechanism allows the realization of a multinode quantum router without the need for complex external controls. By harnessing the bound-state mechanism, we have designed a scalable non-Markovian quantum router that simultaneously solves the challenges of distance limitations, fidelity degradation, and parallel interconnections faced by conventional quantum routers. Our finding establishes an engineered solid-state system as a viable platform for quantum routers and opens an avenue for developing quantum networks.
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (Grants No. 12275109, No. 92576202, and No. 12247101), the Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), and the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799).